Mapuleng deposits R1 500 at the end of every month into an account that earns 12,5% interest per year, compounded monthly. After two years, she stops making these monthly contributions because the interest rate changes to 15% per year, compounded every two months. If no withdrawals or deposits are made for four years the balance in the account will be


1
Expert's answer
2022-03-16T10:20:39-0400

"A\ =P\left(1\ +\ \frac{r}{n}\right)^{n\ast t}"


"A\ =1500\left(1\ +\ \frac{0.125}{12}\right)^{12\ast2}"


"A = \ R1,923.55 \ \ for \ the \ first \ 2 \ years"


"A\ =1500\left(1\ +\ \frac{0.15}{6}\right)^{6\ast2}"


"A = \ R 2,017.33 \ \ \ for \ the \ next \ 2 \ years"


"1923.55 \ + \ 2017.33\ = \ 3940.88"


The balance after 4 years will be "R3,940.88"